find the area of a rectangular playground that is 2700^2 ft the length is 3 times the width
Answer by CharlesG2(834)

find the area of a rectangular playground that is 2700^2 ft the length is 3 times the width
do you mean the playground is 2700 square feet (ft^2) or 2700^2 square feet (ft^2) ?
hmmm didn't you already tell me the area? maybe you want the dimensions?
A area rectangle = L length rectangle * W width rectangle
L = 3W
A = 3W * W = 3W^2
answer 1, if A = 2700 ft^2 = 3W^2
900 = W^2
30 = W then L = 90
answer 2, if A = 2700^2 ft^2 = 3W^2
7290000 = 3W^2
2430000 = W^2
100*sqrt(243) = W
1558.846 approx = W
then L = 4676.537
I think surely you meant for answer 1.